### v1

# New and General Type Meromorphic -forms on Curves

Abstract.
In this article, we study the existence of new and general type meromorphic -forms on curves through explicit construction. Specifically, we have constructed a large family of new and general type meromorphic -forms on elliptic and hyperelliptic curves. We also established a connection to the Hurwitz realization problem of branch cover for the Riemann Sphere, which provides an algorithm to determine whether a -form on (of some restricted class) is new or old.
1. Introduction
Let be an algebraically closed differential field of characteristic zero with trivial derivation (i.e. for all ). Let be an irreducible polynomial such that with function field
Consider a first order algebraic differential equation
where stands for the derivative of Then is a differential field with derivation defined by and for all Let be the pair where is the smooth projective curve with function field and is a meromorphic -form on dual to the linear derivation The -form and the derivation are related as follows: for any Moreover, any such pair with can be thought of as a first order algebraic differential equation over ([7, Lemma 5.2]). As an example, the associated pair for the equation is
To understand the algebraic dependency of the solutions of first order differential equations, the authors in [9] study certain geometric properties of the associated pairs. They classify the pairs (hence first order differential equations) into the following types:
where for some is called of exact type;
where for some and is called of exponential type;
where for some where for with is called of Weierstrass type;
In the rest of the cases, is called of general type.
Furthermore, they define a pair to be new111In [3], the term essential is used in place of new. if does not have a proper pullback, i.e., there exists no pair and a morphism with such that Otherwise, is called old.
In the context of function fields, a pair is of exact type (respectively of exponential type, respectively of Weierstrass type) if an only if the differential field contains an element satisfying (respectively respectively where with ). A pair is new if and only if the differential field is simple, i.e., if is a differential subfield of with then either or
In [9, Theorem 2.1 (a)], it is proved that if a set of nonconstant solutions of a new and of general type equation is algebraically dependent, then there exists such that We would like to remark that in [11] Rosenlicht has proved the following: If any two nonconstant solutions of the equation
are algebraically dependent, then He proves the same for the equation
also. In view of [9], both the equations (1.2) and (1.3) are of general type. Rosenlicht considered these equations to prove that a differentially closed field may admit a proper differentially closed subfield.
Furthermore, the new and of general type equations are βirreducible first order differential equations,β and the nonconstant solutions of these equations are βnew functionsβ in the sense of Nishioka ([8]) and Ummemura ([13]). One cannot express any nonconstant solution of a new and of general type equation iteratively in terms of solutions of linear differential equations, abelian functions, and solutions of any other nonlinear first order differential equations. Also, new and of general type equations do not possess the PainlevΓ© property.
In the publications [3] and [9], the existence of new -forms on a curve 222In the paper, we sometimes say a -form on a curve is type, which will mean the pair is of type. Here, stands for the adjectives: exact, exponential, Weierstrass, general. Also, we say a -form on a curve is new to mean that the pair is new. The same for old. was considered. For Hrushovski and Itai prove the following result ([3, Lemma 2.23]): Let be a -form such that has at least two nonzero residues and that no two distinct nonzero residues of are linearly dependent over then is new. Their proof can be extended to show that is also general type. This is the only known result that can be used to construct new and general type -forms explicitly. Though the equations (1.2) and (1.3) are expected to be new, they cannot be proved or disproved.
For curves of genus only some theoretical result on the existence of new -forms is known. See Lemma 2.18 in [3] for the existence of new -forms on elliptic curves. On a curve of genus it is shown that the set of old forms is contained in a countable union of proper subspaces of ([3, Lemma 2.13]). In [9, Theorem 5.1], Noordman et al. prove that given an effective divisor on a curve of genus a βgenericβ element in is a new -form.
The main purpose of the present paper is to show the existence of new and general type meromorphic -forms on a curve by constructing them explicitly. In particular, we construct a large family of new and general type meromorphic -forms on elliptic and hyperelliptic curves. Our approach also partly answers the following algorithmic question for Decide the place of a given pair (or a differential equation) in the classification of first order differential equations. The main results of the paper are as follows:
Theorem 1.1.
Let be either elliptic or a hyperelliptic curve over Given integers and such that 333For elliptic and hyperelliptic curves, is either zero or an even integer.444In case i.e., -forms with no simple poles, we need and There exists a family of new and general type meromorphic -forms on with number of zeros, number of simple poles, and number of poles of order
Theorem 1.2.
555Theorem 1.2 extends the result [3, Lemma 2.23] of Hrushovski and Itai to curves of arbitrary genus.Let be a curve of genus over and be a meromorphic -form on Assume that has at least poles with nonzero residues and that no subset with distinct elements of the set of nonzero residues is linearly dependent over 666The existence of such a -forms on an arbitrary curve is follows from [6, Proposition 1.15, Page 200].. Then is new and general type.
The proof of these theorems effectively uses the behaviour of zeros, poles and residues of meromorphic -forms under pullback and can be thought of as an application of the Riemann-Hurwitz formula. Theorem 1.2 is sharp in the sense that there are curves of genus and a -form on such that has poles with nonzero residue and is old. For example, let be the smooth hyperelliptic curve given by the affine equation and take
where Then has simple poles, namely and The set of residues can be chosen with no relation except that their sum is zero (which is a property of any -form). Note that is defined in the subfield of The pair is old as it is a pullback of by the natural degree projection
As a corollary, we investigate the existence of a new -form in for an effective divisor on a curve This answers a question proposed in [9, Section 5]. In fact, we consider the existence of a new and general type -form in In general, there may not exist a new -form in For example, let on If then for some The pair is old as it is a proper pullback of via the morphism We prove the following result as a corollary of Theorem 1.2:
Corollary 1.1.
Let where be an effective divisor on a curve of genus
If there exists a family of new and general type -forms in
If there exists a new -form in
We also derive a lower bound of the size of new and general type -forms in for some restricted class of (Theorem 4.10 ). Regarding this, we construct new and general type -forms with a given polar divisor. The result for reads, if with then there exists a vector space consisting only new and general type -forms of dimension at least Here means the greatest integer less than or equal to
We present an algorithm that decides whether a -form on is new or old. Given a pair if it is a proper pullback of another pair via a morphism then we show that the possible branch data of can be determined by Thus, if one can realize certain branch data by a morphism then the type of the -form can be decided. The realization of branch data by a morphism is the classical Hurwitz realization problem for the Riemann sphere. It is still an open problem. But for some particular classes of branch data, the answer to the realization problem is known (see Section 2.1). As a result of that, our algorithm is always applicable to the equations of the form
Consequently, the place of differential equations (1.4) can be decided in the classification. In the end, we prove some explicit criteria for new and of general type equations and show that the equations (1.2) and (1.3) are new and of general type.
Organization of the Paper. In Section 2, we briefly review the Hurwitz realization problem, its solutions, and the existence of meromorphic -forms on curves. Section 3 is about the pullback of pairs and general type -forms. In Section 4, the main results are proven. The algorithm and its applications are displayed in Section 5, with some explicit criteria and examples of new and general type equations at the end.
Acknowledgements. The author would like to thank Varadharaj Ravi Srinivasan and Chetan Balwe for the many helpful discussions, suggestions, and references they provided during the preparation of this article. The author is grateful to IISER Mohali for providing a PhD fellowship.
2. Hurwitz realization problem and existence of meromorphic -forms on curves
2.1. Hurwitz realization problem of branch covering for the Riemann sphere.
Let be a -fold branch covering, or equivalently, a meromorphic function on the Riemann sphere with be its branch points. Then for every there exist points and positive integers such that
β’
for some
β’
and
β’
for every the map on some neighbourhood of with looks like
Thus, a branch covering defines for each of its branch points a set as a partition of The collection is called the branch data of denoted by In this language, the Riemann-Hurwitz formula implies
Definition 2.1.
Let be positive integers. An abstract branch data of degree is a collection where each is a set of positive integers such that for some and satisfy the equation (2.1).
A natural existence problem is the following: for which abstract branch data does there exist a meromorphic function on the Riemann sphere such that In case it exists, call the branch data realizable.
The existence problem for a meromorphic function is a sub-case of the existence problem for branched covering between closed Riemann surfaces, which is posed by Hurwitz ([4]). Many authors studied this problem; for a survey, see [2]. In general, the answer to the above existence problem is negative, i.e., there are non-realizable abstract branch data. For example, the branch data is not realizable. There are some important particular cases when the answer to the existence problem is known:
(Polynomial branch data) A branch data such that there is an for which equivalently and is realizable. A meromorphic function is polynomial if and only if there exists such that and in Hence, the realizability problem is equivalent to the existence of polynomial function. This was proved much earlier in [12].
(Laurent polynomial branch data) A branch data such that there is an for which equivalently and for some is always realizable if for the complete list of non-realizable data is known. The realizability problem is equivalent to the existence of Laurent polynomials. For reference, see [10].
All the branch data of degree are realizable, and for there is only one non-realizable data, namely
2.2. Existence of meromorphic -forms on curves.
In the rest of the paper, will mean a smooth elliptic curve defined by the affine equation will denote a smooth hyperelliptic curve with affine equation The point at infinity will be denoted by Here, we record some known results on the existence of a certain class of meromorphic -forms on curves.
Given a divisor of degree on it is easy to construct a -form with This existence also follows from [6, Problems VI.3]: Let be a divisor on an algebraic curve of genus such that and Then is a canonical divisor. Therefore, for any divisor of degree is a canonical divisor. For arbitrary curves, there are some classes of -forms whose existence is well known. The next proposition implies the existence of -forms on a curve satisfying the assumption of Theorem 1.2.
Proposition 2.1.
[6, Proposition 1.15, Page 200] Given an algebraic curve a finite set of points on and a corresponding subset of there is a meromorphic -form on whose only poles are simple poles at with for each if and only if
Next, we shall describe uniformizer parameters on elliptic and hyperelliptic curves and use them to construct a class of meromorphic -forms.
(1) Elliptic curve: A point on the elliptic curve is called special if where is the inverse of in the elliptic group law. Otherwise, is called ordinary. The elliptic curve has four special points, namely and Let be the uniformizer at For an ordinary point for all and
On define the following class of -forms:
where positive integers, are integers, and are distinct ordinary points of Then
(2) Uniformizer of hyperelliptic curve: An hyperelliptic involution is a morphism defined by for A point is called special if Otherwise, is called ordinary. The curve has special points, namely and Let be the uniformizer at For an ordinary point for all and
On define the following class of -forms:
where positive integers, are integers, and are distinct points of Then
In the class of -forms defined above, if is either zero or an even integer, then is defined in the subfield of Hence is an old form. If is an odd integer, the order of at the special points is either zero or an even integer. Then it follows that given numbers such that is either zero or even number; one can construct a -form on an elliptic or hyperelliptic curve having support at special points with number of zeros, number of simple poles, and number of poles of order Indeed if odd, one can use a special point and number of ordinary points.
3. Pullback of pairs and general type -forms
In this section, we note some observations on the behaviour of zeros, poles, and residues of -forms under pullback. A pair is called pullback of another pair if there is a nonconstant morphism such that Denote it by Ramification index at a point will be denoted by
Lemma 3.1.
Let with Then
Proof.
Let Write where is a uniformization parameter at with We have where is a uniformization parameter at with Then
implies
β
Lemma 3.2.
Let with and where all for all Then the following holds:
for all and
for all divides and
for all divides and Also for any is not a zero of if and only if
if for some then
The same is true for zeros for and also.
If is a pole of with residue , then each is a pole of with residue
Proof.
The proof follows from equation (3.1). β
Proposition 3.1.
If is of exact type, then has a pole of order and all poles of are of order with zero residues.
If is of exponential type, then has a simple pole, and all poles of are simple with residues for some unique and vary with simple poles.
If is of Weierstrass type, then is holomorphic.
If has a pole of order with nonzero residue, then is of general type.
If has a pole of order and a simple pole, then is of general type.
Proof.
By definition, a pair is of exact type (respectively of exponential type, respectively of Weierstrass type) if and only if is pullback of the pair (respectively respectively ). The -form has only one pole of order with zero residues and the -form has only two simple poles with nonzero residues. The -form on is holomorphic. Now the proof of the proposition follows from Lemma 3.2. β
Corollary 3.1.
([5, Remark 7.5])
is of exact type if and only if the partial fraction expression of is of the form
where and is a polynomial over Furthermore, is of exact type and new if and only if for some
is of exponential type if and only if the partial fraction of is of the form
where and are nonzero integers.
Proof.
The proof follows from Proposition 3.1. β
Proposition 3.2.
([9, Proposition 3.3])
If and is of general type then is also of general type.
If is of general type, then has at least one zero.
Proposition 3.3.
Let and be general type. Then where denotes the divisor of zeros of and is the number of zeros of
Proof.
As is general type so is Let be a zero of and By Lemma 3.2, for each is a proper divisor of implies Hence
β
4. Existence of new and general type -forms
4.1. Proof of main theorems and related examples
Let and be two curves with genus and respectively. Let be a morphism with By Riemann-Hurwitz formula,
If then If then implies Therefore, if with and is either a zero or pole of then
For let Note that implies
Lemma 4.1.
Let be a meromorphic -form on a curve of genus with
Let and If for and there is an such that for then is new and general type.
Let and If are prime, there is an such that for and some is greater than equal to then is new and general type.
Proof.
(i) By Proposition 3.1, is of general type . If is old, then there is a morphism and a pair such that with The pair is of general type; hence has a zero, say at (Proposition 3.2). Let for some Then every element in is a zero of and every element in is a simple pole of
We claim that for all If then by equation (4.2), Next, let By Riemann-Hurwitz formula
Then for any pole of
This implies, Also by Lemma 3.2 (ii), divides Since does not have divisor less then equal to Hence for all This proves the claim.
Now for all and implies must contains for some Then by Lemma 3.2 (iv) a contradiction. Therefore, no such morphism and a pair exists, and is new.
(ii) We shall show that there is no morphism and a -form on a curve such that with Let us assume the contrary. If then using the same argument as in the proof of (i), we get for all and the assumptions on lead to a contradiction.
Let We divide the proof into two parts. First, we show that there exists no pair such that has at least one zero and at least two poles of order As has two poles of order by Lemma 3.2 (ii), for each there is a such that Otherwise, for any which contradicts that has two poles of order Let be a zero of By Riemann-Hurwitz formula, for any
Hence for all Then again, using the same argument as in (i), because of the assumptions on we have for some a contradiction.
To complete the proof, we need to show that there is no morphism and -form with with where either has no zero (hence only has a pole of order ) or has one zero and only one pole of order In the first case777This in fact proves that is not exact type. Also, in Lemma 4.1 (ii) cannot be exponential and Weierstrass type, follows from Proposition 3.1., by Lemma 3.2 (iii), Also all poles of will map to the only pole, say of order of Then for and But by assumption, a contradiction. Let be the only pole of order of in the second case. Then all the poles of will map to As are prime, is either or But implies all Again by the same argument, for some a contradiction. Therefore, does not have a proper pullback, and is new and of general type. β
Remark 4.2.
In Lemma 4.1, (i) can be replace by (i) Let and If and for then is new and general type.
Proof of Theorem 1.1.
The proof follow from Lemma 4.1, once we prove that there exists -forms on elliptic and hyperelliptic curves satisfying conditions of Lemma 4.1.
- (i)
Let In this case, -forms satisfying conditions of Lemma 4.1(ii) exists if The last condition ensures that after choosing one of the big enough, one can still choose other zeros of order at least one.
β
Next, we use Lemma 4.1 to produce some easy examples of new and general type -forms on elliptic and hyperelliptic curves.
Examples 4.3.
Let be projective line over
Let and Then Choose and They satisfy the conditions of Lemma 4.1. To get a new and general type -form, we need to choose such that i.e., There are many choices and for each choice we get a -form. For example, Let then the -form is
- (ii)
Examples 4.4.
Examples 4.5.
Let be the hyperelliptic curve of genus given by over
Let Then The following -form on defined by
The pair is new and of general type.
Let Then and Consider the -form on defined by
with Then is new and general type.
Proof of Theorem 1.2.
The -form is general type by 3.1. We show that there exists no pair and a morphism such that with Let be the genus of Then and by the Riemann-Hurwitz formula
Let be the poles of with residues and be the poles of with residues . Then by Lemma 3.2 (v), if for some then where is the ramification index at
We claim that there exists such that in the same fiber for some Let say the claim is true. Then
a contradiction on the assumptions on Hence no such and exists.
Next, we shall prove the claim by contradiction. If the claim is not true, then for each pole of the fiber is a pole of for some unique and Then Also has at least poles with nonzero residue. Otherwise, two or more poles of with nonzero residues will map to a single pole of Then, the residues at the points in the fiber of a pole of will be linearly dependent, a contradiction. By Riemann-Hurwitz formula, the maximum possible number of points such that is ( with ). It follows that has at most poles with nonzero residue; hence so does But then the sum of these nonzero residues is zero, implying that they are dependent. β
Remark 4.6.
If are -form on a curve satisfying the assumptions on Theorem 1.2, hence they are mew and general type -form on Then any linear combination is also a new and general type -form on
4.2. Existence of new and general type -from in
Let be an effective divisor on a curve and In this subsection, we investigate the existence of a new and general type -form in
Proof of Corollary 1.1.
Remark 4.8.
For we have a complete answer on the existence of new and general type -forms in
Proposition 4.1.
Let be an effective divisor on . There exists a new and general type -form in for all except and In the last case, a new - form does exist.
Proof.
Let be an effective divisor on By Corollary 1.1, it is enough to assume We may assume that We will divide the proof in the following cases:
Case (i): Let and A -form on with only simple poles must have at least two poles. If has only two simple poles, then (up to a change of variable), which is exponential type.
Case (ii): Let and There always exists a new and general type -form, as the following example suggests. Let
with It is a general type by Proposition 3.1. To show it is new, let and Since has only one pole of order and But also divides which implies a contradiction.
Case (iii): Let and Let If one of the is greater than the existence follows from Theorem 1.1. Let and New and general type -form exists. Consider the -form
with Then and it is a general type by Proposition 3.1. Now if with then is also general type. Hence has a zero, say at and which implies By equation (3.1), which is not possible.
Let There are no new and general type -forms in If must have a pole at of order such that Let
Since Then the partial fraction of will be
By Corollary 3.1, is exact type. Note that is new. β
Next, we shall calculate a minimum bound of the size of new and general type -form on In general, the subset of new and general type -forms in may not be a vector space (can be seen by Corollary 3.1). However, we shall construct a class of new and general type -forms whose -linear combinations are also new and general type.
Lemma 4.9.
Let with be an effective divisor on Let be a -form having at least one simple pole, at least two poles of order with nonzero residues and polar divisor If there is an such that for and two nonzero residues of at the poles of order are linearly independent over then is new and general type.
Proof.
The -form is general type (Proposition 3.1). We shall show that is new. If not, let with If using the same argument as in Lemma 4.1 (ii), one can show a contradiction. Let First, let has at least two poles of order Then by Lemma 3.2 (ii) for each there exists an such that By Riemann-Hurwitz formula, for any
implies Then and for each there exists such that a contradiction.
Let has only one pole at of order Then inverse image of contains all the poles of order of which will contradict the assumption that has two linearly independent residues at the poles of order β
Now the question is for which effective divisors on elliptic or hyperelliptic curve, one can construct such a -form satisfying the conditions of Lemma 4.9. We have the following answers:
On let where Consider the -form
The polar divisor of is with
On let where all not all zero even positive integer. Consider the -form
where The polar divisor of is with for and and
On let where all and not all zero even positive integer. Consider the -form
where The polar divisor of is with for and with
Theorem 4.10.
Let with be an effective divisor on Let be the third element in the ordered set and Then there exists a vector space consisting only new and general type -forms of dimension at least
Let where all and not all zero, be an effective divisor on the elliptic curve Let be the third element in the ordered set and Then there exists a vector space consisting only new and general type -forms of dimension at least
Let where all and not all zero, be an effective divisor on a hyperelliptic curve Let be the third element in the ordered set and Then there exists a vector space consisting only new and general type -forms of dimension at least
Proof.
As the idea of the proof is similar for all parts of the theorem, we only write the proof of (i). First, construct a -form with polar divisor and - linearly independent residues at Since by Lemma 4.9, is new and general type. For any nonzero is also new and general type. Next, construct another new and general type -form same way as but with polar divisor For any nonzero let Then and the polar divisor of is
Since residue is a -linear map, for and for Then satisfies the conditions of Lemma 4.9. Hence, is new and general type. Observe that and are linearly independent over
Note that in the construction of if one chooses instead of and constructs another -form then all of the above arguments are true. Using these ideas, one can construct linearly independent new and general type -forms in The number of such -forms are at least β
5. Algorithm and Its Applications
5.1. An Algorithm.
Let be a pair. Determining the place of in the classification of first order differential equations is an interesting algorithmic question. We are not aware of an algorithm that answers the question in full generality. There are algorithms that can decide whether an equation is of exact or of exponential type. The first one is an application of the Coates algorithm ([7, Algorithm 2.11]), and the second one follows from the work of Baldassarri and Dwork ([1, Section 6]). Sometimes Proposition 3.1 may be helpful.
We produce an algorithm that decides whether a -form on is new or old. The main idea is that given a -form on if there exists a morphism and a -form on such that then the possibilities of branch data of at the zeros and poles of is determined by the order of zeros and poles of (see Lemma 3.2). All the ramified points of may not come from the set of zeros and poles of But those ramified points outside the zeros and poles of will not take part in finding a suitable Hence, one can find a set of possible abstract branch data for Thatβs where the Hurwitz realization problem comes into play and decides whether there exists a meromorphic function on with given abstract branch data. As we know in Section 2.1, the Hurwitz realization problem for the Riemann sphere is not fully solved. Hence, our algorithm is only applicable to the following -forms on that satisfy one of the following:
the number of zeros of is less than equal to
the number of simple poles of is less than equal to
the number of poles of order of is less than equal to
where denotes the divisor of zeros of and is the number of zeros of
Note that if is exact type, then whether it is new or old can be decided by Corollary 3.1. If is an exponential type, then it is always old. Hence, to decide whether a -form on is new or old, we can restrict ourselves to general type -forms.
Algorithm 5.1.
Β
Input: A general type -form on from the applicable cases.
Output: Decide whether is new or old. If it is old, determine a morphism and -form on such that
1)
Compute all the zeros, poles and residues of
2)
Calculate the possible degrees of a possible morphism by Proposition 3.3. If is general type and has either a single zero or a single pole of order say at then divides In that case, can be excluded.
3)
Fix a possible degree of .
4)
Using Lemma 3.2, calculate the possible ramification index at all the zeros and poles of order The Ramification index at the simple poles is bounded by degree.
5)
From the above data and the Riemann-Hurwitz formula, create a possible abstract branch data, keeping in mind that ramified points of may come from outside the zeros and poles of As is from applicable cases, the abstract branch data will be of the type for which the Hurwitz problem has a solution.
6)
If there is branch data, then calculate the zeros and poles of a possible -form
7)
If there is no branch data or there is branch data with no possible -form, repeat the same process for all possible degrees.
8)
If there is branch data and a -form and the branch data is realizable, then is old. Otherwise, is new.
Β
Consider the differential equations of the form where either or is a Laurent polynomial. The corresponding pair is where One can completely determine its place in the classification of first order differential equations. They canβt be of Weierstrass type. Whether it is of exact, of exponential or of general type can be decided by Corollary 3.1. As or is a Laurent polynomial, can have at most two zeros, two simple poles, or two poles of order Hence, the above algorithm is applicable for this pair.
5.2. Applications
We give a few examples that demonstrate how the algorithm works. For a small number of zeros and poles the algorithm can be performed by hand calculation.
Example 5.1.
The differential equation is new and of general type.
The pair associated to the equation is where
with
By Proposition 3.1, is general type . If then is also of general type. Since has only one zero at , the degree divides The only possible is By the Riemann-Hurwitz formula, Also Note that Consider the following table for calculations:
| Table | ||
For possible branch data are and As is the first case is not possible. is not possible as ramification index at , and cannot be so the other ramification point cannot come from If it comes from outside of then all is They should be in the fiber of the only pole of which is not possible. To get we need two ramified point of ramification index and two unramified point form the set , which is not possible. It cannot come from outside that set because inverse image of poles of are poles of Hence is new.
Example 5.2.
The equation old and of general type.
The pair corresponding to the differential equation is where
By Proposition 3.1, is general type. Since has only one zero of order at and only one pole of order at we have As divides The only possible is We shall check if with
Here For a degree map, only branch data is Hence Then and Hence and and If we assume and then and The equation is old.
Example 5.3.
Consider the differential equation where
The -form is general type by Proposition 3.1 and
If with then by Proposition 3.3, possible βs are and Consider the following table for calculations:
| Table | ||
Observe that and are unramified points and by Lemma 3.2 (ii), maps to a pole of order Also all the can not be and can not be in the same fiber.
Case d=2: The only branch data for a degree map is Then and must contain another zero of As that zero must be an unramified point. This is not possible because
Case d=3: Here must be Otherwise, and is zero of the inverse image must contain either two unramified zeros or one ramified zero with ramification index which is not possible. We shall now look for possible ramification index at and The -form is a general type; hence, must have a zero. That implies both and cannot be Let Then is a zero of and the inverse image must contain either two unramified zeros or one ramified zero with ramification index which is not possible.
There are some explicit criteria for new -forms in a few circumstances. Regarding this, we have the following results:
Proposition 5.1.
Let be a general type meromorphic -form on a curve If has only one zero of order such that is prime, then is new.
Proof.
If is not new, let with Let be the only zero of . Since is a general type, it has a zero at of order By Lemma 3.2, Then and a contradiction. β
Example 5.4.
Proposition 5.2.
Let be a general type meromorphic -form on a curve having only two zeros, one of order and the other of order such that both and are prime, then is new.
Proof.
If is not new, let with Let be two zeros of of order and respectively. As is a general type, it has a zero at of order Three cases could occur: (i) (ii) (iii) In the first and second cases, we have and respectively, which is absurd. In the last case, we get a contradiction. β
Example 5.5.
Consider the equation
where are finite index set and are distinct and Let the set where for some and appears times in the equation. Let be a proper divisor of except and The equation is new and of general type if does not divide all
Proof.
The pair associated to the equation is with
By Proposition 3.1, is general type. Let with As is general type so is Let be a zero of Then with and divides The possible degrees of are divisors of Β except and Also with By Riemann-Hurwitz formula, for all Let be a pole of order of Then by Lemma 3.2, any two poles of in must have same orders. Let be a divisor of i.e. is a possible degree of Each fiber of a pole of of order contains elements and thus divides every a contradiction. Therefore, is new and general type. β